OpenStudy (mathpandas):

How do you solve 13+x/2=3x?

OpenStudy (mathpandas):

Thanks, but I want to know the steps to complete it

OpenStudy (mathmale):

You must group the x-terms on one side of the equation and the constant(s) on the other. I'd suggest you first eliminate the fraction from this problem.

OpenStudy (mathpandas):

If you can, please tell me step by step, thank you, it would help a lot

OpenStudy (mathpandas):

I tried but I got x on both sides

OpenStudy (mathmale):

Multiplying all 3 terms of this equation by 2 would effectively remove fractions from this problem. I can give you feedback ONLY on work that you share with me.

OpenStudy (mathpandas):

Ok

OpenStudy (mathmale):

So, after you've multiplied all three terms by 2, y ou have ... what?

OpenStudy (mathpandas):

You have 13+2=6x?

OpenStudy (mathpandas):

I meant 26+x

OpenStudy (mathpandas):

Help?

OpenStudy (mathmale):

Our answers differ. You have 13+x/2=3x and are to mult. all terms by 2. That gives you 26+x=6x.

OpenStudy (mathmale):

Subtract x from both sides and simplify the result. What do you get?

OpenStudy (mathmale):

Your task here is to group all "x" terms on one side of your equation and all constants on the other.

OpenStudy (mathmale):

Please continue with this problem until you have solved it.